OpenStudy (anonymous):

please help. 2 cars leave from the same point @ the same time, traveling in opposite direction. One car travels 15 mph slower than the other, after 6 hrs they are 630 miles apart. Find the rate of each car.

OpenStudy (anonymous):

let the speed of first=x mph speed of second car=x-15 mph Relative speed v=x+x-15=2x-15 time t=6 hrs distance S=630 miles s=v*t 630=(2x-15)*6 630=12x-90 12x=630+90=720 x=720/12=60 speed of first=60 mph speed of second car=60-15=45 mph

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw